# Exploiting synergies between JWST and cosmic 21-cm observations to uncover star formation in the early Universe

### v1

@@ 2.3.2 Diffuse X-ray background
X-ray emission is another key ingredient of the early Universe, contributing to both the heating of the IGM and the ionization of the neutral hydrogen. Despite its importance, much remains uncertain regarding its properties: soft X-rays ($E\sim 20\,\mathrm{eV}$) in the form of thermal emission from supernovae remnants (Oh, 2001; Venkatesan et al., 2001; Ricotti & Ostriker, 2004) and mini-quasars (Kuhlen et al., 2006; Ciardi et al., 2010), hard X-rays ($E\sim 3\,\mathrm{keV}$) from black hole X-ray binaries (XRBs; Mirabel et al., 2011; Fragos et al., 2013a), and AGN have all been suggested as dominant sources of X-ray emission. The properties of the X-ray background such as its strength, inhomogeneous heating effect (Pritchard & Furlanetto, 2007; Fialkov & Barkana, 2014), and shape of its spectral energy distribution (SED; Fialkov et al., 2014; Fialkov & Barkana, 2014; Pacucci et al., 2014) affect the evolution of the IGM temperature, and thus the 21-cm signal coupled to it. In addition to this, the contribution of ancient X-ray sources to the unresolved X-ray background seen today could act as a useful observable of the early Universe (Fialkov et al., 2017), setting upper limits on their emission efficiency.
@@
Figure 2: UV luminosity functions (UVLFs) in the simulation for a fiducial parameter set: ${V_{c}=4.2\,\mathrm{km}\,\mathrm{s}^{-1}},{\tilde{f}_{\star,\mathrm{II}}=10^{-3}},{M_{0}=5\times 10^{9}\,\mathrm{M}_{\odot}},{\alpha_{\star}=1.0},{\beta_{\star}=0.5}$. Left: UVLF at $z=8$ for a random sample of $\sim 1000$ pixels (grey), averaged across the simulation volume with and without dust effects (black solid and black dash-dotted lines respectively), and compared to a simple analytic model with Sheth-Tormen HMF (black dashed). The suppression at the bright end is a result of dust attenuation and finite simulation volume (purple shaded region), while the suppression at the faint end is due to photoheating feedback during the EoR (orange shaded region). The finite volume and photoheating feedback are simulation features not captured by the analytic model. Right: Redshift evolution of the UVLF from $z=6$ to $z=15$ for the same parameter set.
@@
Fragos et al. (2013b) find XRBs to be dominant at $z\gtrsim 8$. A simple model for their X-ray luminosity, following present-day starburst-like galaxies, assumes proportionality to the local star-formation rate (SFR; e.g. Grimm et al., 2003; Mineo et al., 2012):
$L_{\text{X}}(z)=f_{X}\times\left(3\times 10^{40}{\,\mathrm{erg}\,\mathrm{s}^{-1}}\right)\left(\frac{\text{SFR}}{\,\mathrm{M}_{\odot}\,\mathrm{yr}^{-1}}\right)$ (16)
where $f_{X}$ is the X-ray emission efficiency of sources at high-redshifts, normalized to the prediction from low-metallicity X-ray binaries (Fragos et al., 2013a). The present-day specific intensity from these sources at $z>z_{0}$ can then be calculated using (Pritchard & Furlanetto, 2007):
[…]
@@
usually quoted in units of $\,\mathrm{erg}\,\mathrm{s}^{-1}\,\mathrm{eV}^{-1}\,\mathrm{cm}^{-2}\,\text{sr}^{-1}$. In this equation, $\epsilon_{\rm X}(z^{\prime},E)$ is the specific X-ray emissivity averaged in spherical shells between emission redshift $z^{\prime}$ and present day, $\tau_{\rm X}(z^{\prime},E)$ is the optical depth of X-rays calculated using weighted cross section of hydrogen and helium species (Verner et al., 1996), and $H(z^{\prime})$ is the Hubble parameter at emission. The observed integrated flux in some energy band is then given by:
[…]
@@
in units of $\,\mathrm{erg}\,\mathrm{s}^{-1}\,\mathrm{cm}^{-2}\,\text{sr}^{-1}$. In the case of our simulations, we use $z_{0}=z_{\text{min}}=6$, which means only high-redshift sources are included in the redshift integral. This complements the idea that the X-ray background measurements serve as an absolute upper limit on the contribution from high-$z$ sources. Note that we assume ${f_{X}=}f_{X,\mathrm{II}}=f_{X,\mathrm{III}}$ in the constraints here; the largest contribution to the integral comes from lower redshifts, where Pop II stars likely dominate the star-formation rate.
@@ 2.3.3 Diffuse radio background
$L_{\mathrm{r}}(\nu,z)=f_{r}\times\left(10^{22}\,\text{W}\,\mathrm{s}^{-1}\,\mathrm{Hz}^{-1}\right)\left(\frac{\nu}{150\,\mathrm{MHz}}\right)^{-\alpha_{r}}\left(\frac{\text{SFR}}{\,\mathrm{M}_{\odot}\,\mathrm{yr}^{-1}}\right)$ (19)
where $f_{r}$ is the emission efficiency with respect to the present-day radio galaxies, and the spectral index $\alpha_{r}=0.7$ matches the observed power spectrum of low redshift radio galaxies (Hardcastle et al., 2016; Gürkan et al., 2018). As discussed in the introduction, the excess radio background seen today (Fixsen et al., 2011; Dowell & Taylor, 2018), if non-instrumental, could be due to a variety of exotic sources. The above formalism attributes this, at least in part since the background is an upper limit, to astrophysical origins in galaxies.
To estimate the contribution of high-redshift galaxies at $z>z_{0}$ to the present-day sky-averaged excess radio temperature, we use the equation:
[…]
where $\epsilon_{\rm r}(z^{\prime},\nu)$ is the comoving specific radio emissivity, ${z_{0}=z_{\text{min}}=6}$ as before, and our assumption that ${f_{r}=f_{r,\mathrm{II}}=f_{r,\mathrm{III}}}$. Unlike X-rays, the radio emission is not absorbed as it propagate through space and redshift. We include the effect of line-of-sight fluctuations introduced recently in 21cmSPACE in Sikder et al. (2024a). We also note that we have not included the effect of soft photon heating from the excess background, which can have leading order effects on the matter temperature and thus the 21-cm signal (Cyr et al., 2024), and leave this for future work.
@@ 2.4.1 Analytic model
$\Phi(\mathbf{x},M_{\text{UV}},z)=\dfrac{dn(\mathbf{x},M_{h},z)}{dM_{h}}\left|\dfrac{M_{h}}{M_{\text{UV}}}\right|,$ (21)
with the hybrid HMF $dn/dM_{h}$ discussed in Section 2.1. In order to map $M_{h}\leftrightarrow M_{\text{UV}}$, we assume that the SFR is proportional to the rest-frame UV luminosity of galaxies $L_{\text{UV}}$,
$\dot{M}_{\star}(\mathbf{x},M_{h},z)=\dfrac{M_{\star}(\mathbf{x},M_{h},z)}{t_{\star,\mathrm{II}}H(z)^{-1}}=\kappa_{\text{UV}}\times L_{\text{UV}}(\mathbf{x},M_{h},z)$ (22)
where $M_{\star}$ is the term in brackets in Equation 1, $t_{\star,\mathrm{II}}=0.2$ is the characteristic time for formation of Pop II stars in units of Hubble time, as discussed before in Section 2.1, and $\kappa_{\text{UV}}=1.15\times 10^{-28}\,\mathrm{M}_{\odot}\,\mathrm{yr}^{-1}\,\mathrm{erg}^{-1}\,\mathrm{s}\,\mathrm{Hz}$ is a constant conversion factor. We adopt this value, following Madau & Dickinson (2014) who calculate it assuming continuous star-formation history, a Salpeter IMF in the range 0.10.10.1\,0.1 $\,100\,\mathrm{M}_{\odot}$, and an evolving metallicity of $Z_{\star}=10^{-0.15z}Z_{\odot}$. Note that these assumptions regarding the stellar population are not necessarily justified (e.g. Tacchella et al., 2018; Harikane et al., 2023a) and indeed Madau & Dickinson (2014) caution the metallicity dependence being calibrated to low redshifts only. At high redshifts, the metallicity could be drastically lower, with a different IMF leading to a much larger $\kappa_{\text{UV}}$. We use the value above for consistency with other works, but briefly explore its impact on the UVLF, together with those of $\delta_{\text{max}}$ and $t_{\star,\mathrm{II}}$, in Appendix A.
@@
Although Equation 21 is an analytic prescription, we stress that it is applied at a pixel level ($3\,\mathrm{cMpc}$ scale) before averaging over the simulation volume. In Figure 2, we compare our semi-numerical UVLF calculated from the simulation volume with a simple analytic model derived using the Sheth-Tormen HMF (calculated independently). As expected, they are in good agreement but the simulation-based approach has two benefits: (i) we can capture the ‘extreme statistics’ of rare overdensities in finite simulation volume, and (ii) each pixel has a unique $M_{\text{crit}}$ from photoheating feedback during the EoR, giving rise to a natural suppresion of faint galaxies. The model can also be easily extended to simulate larger volumes and study large-scale overdensity biases, or include spatially varying metallicities which would alleviate the need for a constant $\kappa_{\text{UV}}$. We leave such investigations for future work.
@@ 2.4.2 Dust correction
Physically, this leads to a suppression of the brightest galaxies at redshifts $z\lesssim 10$, but largely does not affect galaxies at $z\gtrsim 10$. This is because galaxies are generally not large enough to reach the high magnitudes where dust effects becoming important, and there is an ‘expected’ absence of dust in this calibration. A potential mechanism for lack of dust in galaxies at high-$z$ (the so-called ‘blue monsters’) is by radiation driven outflows (Ziparo et al., 2023; Fiore et al., 2023; Ferrara et al., 2024).
@@ 2.5 Simulation parameters
$\alpha_{\star}$ Uniform 00 $2$
$\beta_{\star}$ Uniform 00 $5$
@@
SARAS $\sigma_{\text{S3}}$ Log-Uniform $0.01$ K $1$ K
@@
As the early Universe remains uncertain, the cosmological and astrophysical processes described in Section 2.1 can be turned on/off or scaled through various parameters in 21cmSPACE. In this work, we focus on exploring the effect of different star-formation models. Hence, our free parameter space consists of eight parameters: $\big{\{}V_{c},\tilde{f}_{\star,\mathrm{II}},M_{0},\alpha_{\star},\beta_{\star},\tau,f_{X},f_{r}\big{\}}$. The method of obtaining constraints on these parameters using Bayesian analysis closely follows that of previous works (Gessey-Jones et al., 2023; Pochinda et al., 2024; Bevins et al., 2024), but with the updated SFE and inclusion of UVLF observable. We briefly list the parameters and their prior ranges below.
  • $V_{c}$ – Minimum circular velocity of halos for star-formation in the absence of feedback. As described in Section 2.2, the feedback processes increase the $M_{\text{crit}}$ value for each pixel, but $V_{c}$ sets the minimum threshold for star-formation in the absence of feedback processes. We vary it in the range $V_{c}\in\left[4.2,100\right]\,\mathrm{km}\,\mathrm{s}^{-1}$, corresponding to $M_{\text{crit}}\in\left[3\times 10^{6},4\times 10^{10}\right]\,\mathrm{M}_{\odot}$ at $z=6$ (see Equation 3, in the absence of $f_{v_{\text{bc}}}$ and $f_{\text{LW}}$), and $T_{\text{vir}}\in\left[1.3\times 10^{3},7.4\times 10^{5}\right]\,\text{K}$ (see Equation 5).
@@
  • $\alpha_{\star}$ – The power-law index of the high mass end of the SFE, varied in the range $\alpha_{\star}\in\left[0,2\right]$. We choose positive values to ‘boost‘ the SFE at the high mass end in agreement with observations (e.g. Behroozi & Silk, 2015; Behroozi et al., 2019). A value of 00 corresponds to a flat SFE in the atomic cooling regime, while a value of $2$ corresponds to a steep power-law (i.e., efficiency rapidly increasing with halo mass).
@@
  • $\beta_{\star}$ – Redshift evolution of $M_{\text{high}}$, varied in the range $\beta_{\star}\in\left[0,5\right]$. Higher values of $\beta_{\star}$ correspond to a faster evolution of the SFE with redshift, while a value of 00 corresponds to no evolution (i.e., redshift-independence). This parameter is introduced to capture an effective boost in SFE in the early Universe at small halo masses, in order to explain the observed high-$z$ galaxy populations from JWST.
@@
  • $\tau$ – In 21cmSPACE, the effective ionization efficiency of galaxies is captured by $\xi_{\text{ion}}$ is as defined in Equation 2. Since the optical depth of CMB photons to reionization $\tau$ monotonically scales with $\xi_{\text{ion}}$ when $V_{c},f_{\star,\mathrm{II}}$ and other parameters are fixed (Cohen et al., 2020), we use $\tau$ instead of $\xi_{\text{ion}}$ in our analysis as the free parameter. We adopt a uniform prior of $3\sigma$ around the measured Planck 2018 value of $\tau=0.054\pm 0.07$ (Planck Collaboration et al., 2020), i.e. $\tau\in\left[0.033,0.075\right]$. By indirectly using this reionization era constraint, we ensure that no unphysical reionization histories are realized in our models. We do not explore more complex parameterizations here, instead focusing on the SFE, due to uncertainties in the galaxy ionizing emissivities, escape fraction of ionizing sources, and IGM clumping which contribute to $\zeta_{\text{ion}}$ (e.g Mitra & Chatterjee, 2023; Davies et al., 2024; Asthana et al., 2025).
@@
  • $f_{X}$ – The X-ray emission efficiency of high-redshift galaxies, relative to present-day starburst-like galaxies as defined in Equation 16. We vary this in the wide range $f_{X}\in\left[10^{-3},10^{3}\right]$. Since this parameter is degenerate with the SFR of galaxies (and thus the SFE), the wide range aptly allows for extreme parameter combinations such as very low SFE and very high $f_{X}$.
@@
  • $f_{r}$ – The radio emission efficiency of high-redshift galaxies, relative to present-day galaxies as defined in Equation 19. We vary this parameter in the wide $f_{r}\in\left[10^{-1},10^{5}\right]$, chosen to be wide for the same reason as $f_{X}$.
The above described parameter priors are summarized in Table 1. We include radio and X-ray efficiencies as free parameters due to their impact on the 21-cm signal, and degeneracy with star-formation rates in determining the present-day diffuse radio and X-ray backgrounds which we include our constraints.
@@
With the parameters described above, we run a suite of 30,000 simulations across the eight-dimensional astrophysical parameter space. Each simulation takes a couple of hours to run, and the entire suite takes a few weeks to complete on the DiRAC high-performance computing facility COSMA8.666See https://cosma.readthedocs.io/en/latest/cosma8 for more details on CPU specifications.

### v2

@@ 2.3.2 Diffuse X-ray background
X-ray emissions are another key ingredient of the early Universe, contributing to both the heating of the IGM and the ionization of the neutral hydrogen. Despite its importance, much remains uncertain regarding its properties: soft X-rays ($E\sim 20\,\mathrm{eV}$) in the form of thermal emission from supernovae remnants (Oh, 2001; Venkatesan et al., 2001; Ricotti & Ostriker, 2004) and mini-quasars (Kuhlen et al., 2006; Ciardi et al., 2010), hard X-rays ($E\sim 3\,\mathrm{keV}$) from black hole X-ray binaries (XRBs; Mirabel et al., 2011; Fragos et al., 2013a), and AGN have all been suggested as dominant sources of X-ray emission. The properties of the X-ray background such as its strength, inhomogeneous heating effect (Pritchard & Furlanetto, 2007; Fialkov & Barkana, 2014), and shape of its spectral energy distribution (SED; Fialkov et al., 2014; Fialkov & Barkana, 2014; Pacucci et al., 2014) affect the evolution of the IGM temperature, and thus the 21-cm signal coupled to it. In addition to this, the contribution of ancient X-ray sources to the unresolved X-ray background seen today could act as a useful observable of the early Universe (Fialkov et al., 2017), setting upper limits on their emission efficiency.
@@
Figure 2: UV luminosity functions (UVLFs) in the simulation for a fiducial parameter set: ${V_{c}=4.2\,\mathrm{km}\,\mathrm{s}^{-1}},{\tilde{f}_{\star,\mathrm{II}}=10^{-3}},{M_{0}=5\times 10^{9}\,\mathrm{M}_{\odot}},{\alpha_{\star}=1.0},{\beta_{\star}=0.5}$. Left: UVLF at $z=8$ for a random sample of $\sim 1000$ pixels (grey), averaged across the simulation volume with and without dust effects (black solid and black dash-dotted lines respectively). The suppression at the bright end is a result of dust attenuation and a finite simulation volume (left, purple shaded region), while the suppression at the faint end is due to photoheating feedback during the EoR (right, orange shaded region). The finite volume and photoheating feedback are simulation features not captured by simple analytic models. Right: Redshift evolution of the UVLF from $z=6$ to $z=15$ for the same parameter set.
@@
XRBs may be a dominant source of X-ray emissions at $z\gtrsim 6$, in particular HMXBs (HMXBs; Fragos et al., 2013b; Lehmer et al., 2016). A simple model for their X-ray luminosity, following local starburst-like galaxies, assumes proportionality to the halo star-formation rate (SFR; e.g. Grimm et al., 2003; Lehmer et al., 2010; Mineo et al., 2012, 2014):
$\dfrac{L_{X}}{\mathrm{SFR}}=f_{X}\times\left(3\times 10^{40}{\,\mathrm{erg}\,\mathrm{s}^{-1}}\,\mathrm{M}_{\odot}^{-1}\,\mathrm{yr}\right)$ (16)
where $f_{X}$ is the X-ray emission efficiency of sources at high-redshifts, normalized to the theoretical prediction for low metallicity HMXBs (Fragos et al., 2013a, b; Lehmer et al., 2021) as explained in Fialkov et al. (2017). The comoving X-ray emissivity throughout the simulation can then be calculated as $\epsilon_{X}(\mathbf{x},E,z)=L_{X}/\text{SFR}\times\dot{\rho}_{\star,\mathrm{II}+\mathrm{III}}(\mathbf{x},z)\times\hat{\epsilon}_{X}(E)$} where $\dot{\rho}_{\star,\mathrm{II}+\mathrm{III}}$ is the total Pop II and Pop III SFRD, and $\hat{\epsilon}_{X}(E)$ is the normalized X-ray SED in $\mathrm{eV}^{-1}$ (the mean model from Fragos et al., 2013b, at $z=15.34$ in the energy range $0.2-95\,\mathrm{keV}$, including interstellar absorption). The present-day angle-averaged specific intensity from these sources at $z>z_{0}$ is thus (e.g. Pritchard & Furlanetto, 2007):
$J_{X}(E,z=0)=\frac{1}{4\pi}\int_{z_{0}}^{\infty}\epsilon_{X}(E^{\prime},z^{\prime})e^{-\tau_{X}(E^{\prime},z^{\prime})}\bigg{|}\dfrac{cdt}{dz^{\prime}}\bigg{|}dz^{\prime},$ (17)
@@
usually quoted in units of $\mathrm{erg}\,\mathrm{s}^{-1}\,\mathrm{eV}^{-1}\,\mathrm{cm}^{-2}\,\text{sr}^{-1}$. In this equation, $E^{\prime}=E(1+z^{\prime})$ is the energy at the emission redshift $z^{\prime}$, $\epsilon_{X}(E^{\prime},z^{\prime})$ is averaged across the simulation box [^6] There is no need for shell radiative transfer here since we’re interested in the sky-averaged diffuse background., $\tau_{X}(E^{\prime},z^{\prime})$ is the optical depth of X-rays calculated using the weighted cross section of hydrogen and helium species (Verner et al., 1996), and $H(z^{\prime})$ is the Hubble parameter at emission. The observed integrated flux in some energy band is then given by:
$S_{X}\left(E_{\text{band}}\right)=\int_{E_{\text{band}}}J_{X}(E,z=0)dE$ (18)
@@
in units of $\mathrm{erg}\,\mathrm{s}^{-1}\,\mathrm{cm}^{-2}\,\text{sr}^{-1}$. In the case of our simulations, we use $z_{0}=z_{\text{min}}=6$, which means that only high-redshift sources are included in the redshift integral. This complements the idea that the X-ray background measurements serve as an absolute upper limit on the contribution from high-$z$ sources. Note that we assume ${f_{X}=}f_{X,\mathrm{II}}=f_{X,\mathrm{III}}$ in the constraints here; the largest contribution to the integral comes from lower redshifts, where Pop II stars dominate the star-formation rate.
@@ 2.3.3 Diffuse radio background
$\dfrac{L_{r}(\nu)}{\text{SFR}}=f_{r}\times\left(10^{22}\,\text{W}\,\mathrm{s}^{-1}\,\mathrm{Hz}^{-1}\,\mathrm{M}_{\odot}^{-1}\,\mathrm{yr}\right)\left(\frac{\nu}{150\,\mathrm{MHz}}\right)^{-\alpha_{r}}$ (19)
where $f_{r}$ is the emission efficiency with respect to the present-day radio galaxies, and the spectral index $\alpha_{r}=0.7$ matches the observed power spectrum of synchrotron emission from low redshift radio galaxies (Hardcastle et al., 2016; Gürkan et al., 2018). As discussed in the introduction, the excess radio background seen today (Fixsen et al., 2011; Dowell & Taylor, 2018), if non-Galatic, could be due to a variety of exotic sources. The above formalism attributes this, at least in part since the background is an upper limit, to astrophysical origins in galaxies.
To estimate the contribution of high-redshift galaxies at $z>z_{0}$ to the present-day sky-averaged excess radio temperature, we use the equation (e.g. Ewall-Wice et al., 2020; Reis et al., 2020):
$T_{r}(\nu,z=0)=T_{\text{CMB}}+\frac{c^{2}}{2k_{\rm B}\nu^{2}}\dfrac{1}{4\pi}\int_{z_{0}}^{\infty}\epsilon_{r}(\nu^{\prime},z^{\prime})\bigg{|}\dfrac{cdt}{dz^{\prime}}\bigg{|}dz^{\prime},$ (20)
where $\nu^{\prime}=\nu(1+z^{\prime})$ is the frequency at emission redshift $z^{\prime}$, $\epsilon_{r}(\nu,z^{\prime})$ is the comoving radio emissivity averaged across the simulation volume, $z_{0}=z_{\text{min}}=6$} as before, and we assume that $f_{r}=f_{r,\mathrm{II}}=f_{r,\mathrm{III}}$}. Unlike X-rays, the radio emission is not absorbed as it propagate through space and time. We include the effect of line-of-sight fluctuations introduced recently in 21cmSPACE in Sikder et al. (2024a). We also note that we have not included the effect of soft photon heating from the excess background, which can have leading order effects on the matter temperature and thus the 21-cm signal (Cyr et al., 2024), and leave this for future work.
@@ 2.4.1 Analytic model
$\Phi(\mathbf{x},M_{\text{UV}},z)=\big{(}1+\delta(\mathbf{x})\big{)}\dfrac{dn(\mathbf{x},M_{h},z)}{dM_{h}}\left|\dfrac{dM_{h}}{dM_{\text{UV}}}\right|,$ (21)
where $\Phi$ is the number density of galaxies at UV magnitude $M_{\text{UV}}$ (defined later in Equation 23), $dn/dM_{h}$ is the hybrid HMF discussed in Section 2.1, and the $(1+\delta)$ accounts for the conversion of the HMF from Lagrangian to Eulerian space as in Equation 1 (e.g. Tseliakhovich et al., 2011). In order to map $M_{h}\leftrightarrow M_{\text{UV}}$, we assume that the SFR is proportional to the rest-frame UV luminosity of galaxies $L_{\text{UV}}$,
$\dot{M}_{\star,\mathrm{II}}(\mathbf{x},M_{h},z)=\dfrac{M_{\star,\mathrm{II}}(\mathbf{x},M_{h},z)}{t_{\star,\mathrm{II}}H(z)^{-1}}=\kappa_{\text{UV}}\times L_{\text{UV}}(\mathbf{x},M_{h},z)$ (22)
where $M_{\star,\mathrm{II}}$ is the term in parentheses in the integrand of Equation 1, $t_{\star,\mathrm{II}}=0.2$ is the characteristic time for formation of Pop II stars in units of the Hubble time, as discussed before in Section 2.1, and $\kappa_{\text{UV}}=1.15\times 10^{-28}\,\mathrm{M}_{\odot}\,\mathrm{yr}^{-1}\,\mathrm{erg}^{-1}\,\mathrm{s}\,\mathrm{Hz}$ is a constant conversion factor. We adopt this value, following Madau & Dickinson (2014) who calculate it assuming a continuous star-formation history, a Salpeter IMF in the range 0.10.1\, $100\,\mathrm{M}_{\odot}$, and an evolving metallicity of $Z_{\star}=10^{-0.15z}Z_{\odot}$. Note that these assumptions regarding the stellar population are not necessarily justified (e.g. Tacchella et al., 2018; Harikane et al., 2023a) and indeed Madau & Dickinson (2014) caution about the metallicity dependence being calibrated to low redshifts only. At high redshifts, the metallicity could be drastically lower, with a different IMF leading to a much larger $\kappa_{\text{UV}}$. We use the value above for consistency with the rest of literature, but briefly explore its impact on the UVLF, together with those of $\delta_{\text{max}}$ and $t_{\star,\mathrm{II}}$, in Appendix A. We find that the difference between assuming a fixed Chabrier or Salpeter IMF is negligible compared to modelling uncertainties, although noting that more extreme IMFs like top-heavy Pop III stars that are not explored here could generate larger differences (e.g. Inayoshi et al., 2022).
@@
Although Equation 21 is an analytic prescription, we stress that it is applied at a pixel level ($3\,\mathrm{cMpc}$ scale) before averaging over the simulation volume as shown in Figure 2. This simulation-based approach has two benefits over simple analytic models: (i) we can capture the ‘extreme statistics’ of rare overdensities in finite simulation volume, and (ii) each pixel has a unique $M_{\text{crit}}$ from photoheating feedback during the EoR, giving rise to a natural suppresion of faint galaxies. The model can also be easily extended to simulate larger volumes and study large-scale overdensity biases, or include spatially varying metallicities which would alleviate the need for a constant $\kappa_{\text{UV}}$. We leave such investigations for future work.
@@ 2.4.2 Dust correction
Physically, this leads to a suppression of the brightest galaxies at redshifts $z\lesssim 10$, but does not significantly affect galaxies at $z\gtrsim 10$. This is because galaxies are generally not large enough to reach the high magnitudes where dust effects become important, and there is an ‘expected’ absence of dust in this calibration. A potential mechanism for lack of dust in galaxies at high-$z$ (the so-called ‘blue monsters’) is by radiation driven outflows (Ziparo et al., 2023; Fiore et al., 2023; Ferrara et al., 2025).
@@ 2.5 Simulation parameters
$\alpha_{\star}$ Uniform 0 22
$\beta_{\star}$ Uniform 0 55
@@
SARAS $\sigma_{\text{S3}}$ Log-Uniform $0.01$ K 11 K
@@
As the early Universe remains uncertain, the cosmological and astrophysical processes described in Section 2.1 can be turned on/off or scaled through various parameters in 21cmSPACE. In this work, we focus on exploring the effect of different star-formation models. Hence, our free parameter space consists of eight parameters: $\big{\{}V_{c},\tilde{f}_{\star,\mathrm{II}},M_{0},\alpha_{\star},\beta_{\star},\tau,f_{X},f_{r}\big{\}}$. The method of obtaining constraints on these parameters using Bayesian analysis closely follows that of previous works (Gessey-Jones et al., 2023; Pochinda et al., 2024; Bevins et al., 2024), but with the updated SFE and inclusion of the UVLF observable. We briefly list the parameters and their prior ranges below.
  • $V_{c}$ – Minimum circular velocity for star-formation in halos in the absence of feedback. As described in Section 2.2, the feedback processes increase the $M_{\text{crit}}$ value for each pixel, but $V_{c}$ sets the minimum threshold for star-formation in the absence of feedback processes ($M_{\text{vir}}$). We vary it in the range $V_{c}\in\left[4.2,100\right]\,\mathrm{km}\,\mathrm{s}^{-1}$, corresponding to $M_{\text{vir}}\in\left[3\times 10^{6},4\times 10^{10}\right]\,\mathrm{M}_{\odot}$ at $z=6$ (see Equation 3), and $T_{\text{vir}}\in\left[6.35\times 10^{2},3.6\times 10^{5}\right]\,\text{K}$ (see Equation 6).
@@
  • $\alpha_{\star}$ – The power-law index of the high mass end of the SFE, varied in the range $\alpha_{\star}\in\left[0,2\right]$. We choose positive values to ‘boost‘ the SFE at the high mass end in agreement with observations (e.g. Behroozi & Silk, 2015; Behroozi et al., 2019). A value of 0 corresponds to a flat SFE in the atomic cooling regime, while a value of $2$ corresponds to a steep power-law (i.e., efficiency rapidly increasing with halo mass).
@@
  • $\beta_{\star}$ – Redshift evolution of $M_{\text{high}}$, varied in the range $\beta_{\star}\in\left[0,5\right]$}. Higher values of $\beta_{\star}$ correspond to a faster evolution of the SFE with redshift, while a value of 0 corresponds to no evolution (i.e., redshift-independence). This parameter is introduced to capture an effective boost in SFE in the early Universe for small halo masses, in order to explain the observed high-$z$ galaxy populations from JWST.
@@
  • $\tau$ – In 21cmSPACE, the effective ionization efficiency of galaxies is captured by $\xi_{\text{ion}}$ as defined in Equation 2. Since the optical depth of CMB photons to reionization $\tau$ monotonically scales with $\xi_{\text{ion}}$ when $V_{c},f_{\star,\mathrm{II}}$ and other parameters are fixed (Cohen et al., 2020), we use $\tau$ instead of $\xi_{\text{ion}}$ in our analysis as the free parameter. We adopt a uniform prior of $3\sigma$ around the measured Planck 2018 value of $\tau=0.054\pm 0.07$ (Planck Collaboration et al., 2020), i.e. $\tau\in\left[0.033,0.075\right]$ [^7] For late reionization scenarios, beyond the end of our simulation ($z\lesssim 6$), we calculate the $\tau$ by extrapolating the ionization history as a function of redshift in log-log space.. By indirectly using this reionization era constraint, we ensure that no unphysical reionization histories are realized in our models. We do not explore more complex parameterizations here, instead focusing on the SFE, due to uncertainties in the galaxy ionizing emissivities, escape fraction of ionizing sources, and IGM clumping which contribute to $\zeta_{\text{ion}}$ (e.g Mitra & Chatterjee, 2023; Davies et al., 2024; Asthana et al., 2025).
@@
  • $f_{X}$ – The X-ray emission efficiency of high-redshift galaxies, relative to the theoretical predictions of low metallicity HMXBs of Fragos et al. (2013b), as defined in Equation 16. We vary this in the wide range $f_{X}\in\left[10^{-3},10^{3}\right]$. Since this parameter is degenerate with the SFR of galaxies (and thus the SFE), the wide range aptly allows for extreme parameter combinations such as very low SFE and very high $f_{X}$.
@@
  • $f_{r}$ – The radio emission efficiency of high-redshift galaxies, relative to present-day galaxies as defined in Equation 19. We vary this parameter in the range $f_{r}\in\left[10^{-1},10^{5}\right]$, chosen to be wide for the same reason as $f_{X}$.
@@
The above described parameter priors are summarized in Table 1. Using these, we run a suite of 30,000 simulations across the eight-dimensional astrophysical parameter space. Each simulation takes a couple of hours to run, and the entire suite takes a few weeks to complete on the DiRAC high-performance computing facility COSMA8.888See https://cosma.readthedocs.io/en/latest/cosma8 for more details on CPU specifications.